# A Vision-Based Shared-Control Teleoperation Scheme for Controlling the Robotic Arm of a Four-Legged Robot

### v1

@@ I Introduction
Recent advances in legged locomotion and mobile manipulation have significantly expanded robotic capabilities in different applications, including challenging environments. In such scenarios, robotic systems can be employed to perform telemanipulation where direct human presence is not feasible, being employed in tasks ranging from maintenance of industrial systems [1, 2] to search and rescue response in remote, hazardous environments [3, 4]. The integration of robotic arms has expanded the applicability of quadruped robots in areas such as disaster response [5]. This progress allows robots to move around and interact with the world in useful and practical ways to both navigate and physically interact within human-centric, high-risk environments [6].
@@
Human operators frequently manage the control of such robotic systems. Due to the complexity of these remote interactions and frameworks, a high level of training and expertise is required. Despite their sophistication, these robots often rely on classical input devices like joysticks or gamepads for teleoperation. While familiar, these interfaces present steep learning curves and high cognitive loads, especially for non-experts controlling high-DoF manipulators on mobile platforms [7]. Managing a manipulator’s DoF motion is hard to use with joysticks [8], often worsened by poor situational awareness from inadequate environmental perception [6]. Operators often struggle to perceive 3D environments, plan collision-free trajectories, and execute fine motor commands under limited feedback, leading to reduced precision and fatigue [9].
@@
To enhance intuitiveness, human-in-the-loop interfaces integrate human cognitive capabilities with robotic execution. Wearable sensors, like IMU-based motion capture systems or Electromyography (EMG) armbands, can map human movements to the robot [10, 11]. However, wearable devices face deployment drawbacks, such as time-consuming calibration and intrusive physical contact. EMG systems are particularly prone to issues like motion artifacts, fatigue, electrode shifting, impedance variations, and cross-talk between adjacent muscles [12, 13].
@@
Camera-based interfaces offer a non-invasive, accessible alternative. Such systems eliminate the need for physical instrumentation and offer a cost-effective, hands-free setup. However, such interfaces also present implementation challenges. Monocular RGB cameras lack inherent depth information, complicating 3D pose estimation. The absence of direct depth information in monocular systems results in significantly reduced accuracy compared to RGB-D approaches, with studies showing that monocular depth estimation suffers from scale ambiguity [14, 15, 16] and drift over time, particularly affecting precision in dynamic manipulation tasks [17, 18]. Moreover, many vision solutions are limited to specific gestures for discrete actions or use complex pipelines that hinder real-time responsiveness and generalization [19, 20]. RGB-D cameras, providing direct depth measurement, simplify 3D pose estimation for manipulator control and demonstrate superior accuracy compared to monocular approaches [21, 15].
@@
In order to improve the robustness and applicability of human-machine interfaces, intelligent control systems can assist by partially automating task execution, reducing the overall complexity of the task, and lowering the user’s effort. Semi-autonomous systems are presented as an alternative to complex and counterintuitive manual systems, combining decoded user intentions with autonomous control modules. This work introduces a vision-based teleoperation shared-control framework designed to overcome the limitations outlined above, providing intuitive, real-time control of a quadruped’s manipulator. Our system uses a depth camera for robust 3D perception and a streamlined hand pose estimation pipeline. The proposed framework tracks the operator’s wrist position for continuous end-effector control and recognizes distinct hand gestures for discrete commands, enabling natural, robust, and responsive arm teleoperation.
@@
The interface allows easy switching between manual and autonomous modes through finger counting, offering a simple and low-effort way to share control. In autonomous mode, the robot initiates a grasp when a closed-hand gesture is detected over a detected object. The system identifies graspable items from a predefined list, picks the one with the highest confidence and closest proximity, and provides its location to help the robot find and pick it. This combination of depth vision, simplified hand tracking, object recognition, and intuitive mode switching aims for a holistic, practical system. Importantly, it requires neither wearable devices nor traditional interfaces, addressing cost and setup barriers and offering a low-cost, accessible alternative for complex telemanipulation, especially where other systems are impractical. Experiments demonstrated a favorable balance of intuitiveness, responsiveness, and practicality for a series of pick-and-place tasks, as shown in Fig. 1.
@@
The rest of the paper is organized as follows: Section II explains how the framework was implemented and deployed, Section III presents the apparatus used and the experiments performed to evaluate the efficiency of the framework, alongside the experimental results of this study. Section IV concludes the paper.
@@ II-A Camera-Based Motion Capture System
Figure 2: Overview of the proposed control pipeline, illustrating the flow from motion capture to robot actuation. A manual teleoperation module allows the user to control the robot in an intuitive manner using an external camera. A semi-autonomous module, capable of autonomously grasping an object, can be triggered by the user by closing their hands.
@@ II-B Control Framework
The Spot robotic arm is controlled using the Boston Dynamics Software Development Kit (SDK) [26], ROS-based interfaces, and the MoveIt motion planning framework [27], enabling precise end-effector pose control. Motion commands are generated based on the estimated 3D wrist pose of the user, $p_{\text{marker}}^{\text{wrist}}$, acquired through the camera-based motion capture system detailed in Section II-A. These commands, representing the desired end-effector pose for the robot, are transmitted to the Spot SDK via ROS, enabling real-time control of the arm. Gripper actions, such as opening and closing for object manipulation, are similarly triggered by discrete gesture-based inputs recognized by an independent gesture recognition module based on hand landmarks from MediaPipe. This shared-control architecture enables intuitive teleoperation while ensuring robust and safe object manipulation. The overall control pipeline is illustrated in Fig. 2. This hybrid architecture exploits the advantages of both the proprietary Boston Dynamics SDK, providing robust, low-level access to Spot’s hardware, and the flexibility and extensibility of the open-source ROS ecosystem. ROS, together with MoveIt, manages bidirectional command and state translation, acting as the integration point for perception, planning, and actuation.
@@ II-B4 Implementation Details
The control loop, encompassing motion capture data processing, command generation, and communication with the robot, is developed for real-time performance. Perception tasks (object detection, if active) and motion planning are executed in parallel where feasible. The code for integrating user input, MoveIt functionalities, and dispatching commands to Spot, alongside Docker files, are publicly available in our repository (see footnote in Section II-A). Debugging overlays are provided for visualization of all perception and control stages, aiding parameter tuning and troubleshooting.
@@ III-B Simulation Experiments
Figure 5: Decoded position of the user’s wrist using the pose estimation of the wrist and using the Optitrack system, and the translated robot’s end-effector position captured using the Optitrack system.
@@ III-D Precision analysis
In order to perform a more comprehensive analysis of the proposed framework, we attached reflexive markers to the user’s wrist and robot’s end-effector to capture its position using an Optitrack system comprised of six Primex, 41 cameras and used this as ground truth to perform an analysis of the pose estimation’s positional error, as the Optitrack system can achieve 3D accuracy of up to +/- 0.10 mm. The results are shown in Fig. 5. The positional error between the user’s wrist from the pose estimation and the actual robot’s end effector is shown in Fig. 6, alongside the user’s wrist velocity. A mean positional error of 0.07 m was achieved, and it can be seen that the error is higher when the wrist velocity is higher. This is expected, as the robot’s end effector takes some time to reach the goal position, leading to this increased error.
@@ IV Conclusion
This paper presented an intuitive and assistive vision-based telemanipulation framework for the control of a robotic arm system of a four-legged robot. An external camera is used to track the operator’s arm movement, offering intuitive control of the robotic system. The framework enables the user to control the end-effector position and orientation, where the position is obtained by tracking the user’s wrist position, and orientation tracking is achieved by estimating the pose of the operator’s hand. The proposed system also enables collision detection, ensuring safe operation. The framework was experimentally validated in a telemanipulation task in which two users successfully performed pick-and-place tasks in real-time. Future work will focus on implementing a path planner in the framework so that obstacle avoidance can be achieved. Moreover, we also plan to implement a potential fields scheme [29] in order to improve the robustness of our proposed method through a shared-control approach.

### v2

@@ I Introduction
Recent advances in legged locomotion and mobile manipulation have significantly expanded the capabilities of robots in various applications, including those in challenging environments. In such scenarios, robotic systems can be employed to perform telemanipulation where direct human presence is not feasible, being employed in tasks ranging from maintenance of industrial systems [1, 2] to search and rescue response in remote, hazardous environments [3, 4]. The integration of robotic arms has expanded the applicability of quadruped robots in areas such as disaster response [5]. This progress enables robots to move around and interact with the world in useful and practical ways, both navigating and physically interacting within human-centric, high-risk environments [6].
@@
Human operators frequently manage the control of such robotic systems. Due to the complexity of these remote interactions and frameworks, a high level of training and expertise is required. Despite their sophistication, these robots often rely on classic input devices, such as joysticks or gamepads, for teleoperation. While familiar, these interfaces present steep learning curves and high cognitive loads, especially for non-experts controlling high-DoF manipulators on mobile platforms [7]. Managing a manipulator’s DoF motion is hard to use with joysticks [8], often worsened by poor situational awareness from inadequate environmental perception [6]. Operators often struggle to perceive 3D environments, plan collision-free trajectories, and execute fine motor commands under limited feedback, leading to reduced precision and fatigue [9].
@@
To enhance intuitiveness, human-in-the-loop interfaces integrate human cognitive capabilities with robotic execution. Wearable sensors, like IMU-based motion capture systems or Electromyography (EMG) armbands, can map human movements to the robot [10, 11]. However, wearable devices face deployment challenges, such as time-consuming calibration and the need for intrusive physical contact. EMG systems are particularly prone to issues like motion artifacts, fatigue, electrode shifting, impedance variations, and cross-talk between adjacent muscles [12, 13].
@@
Camera-based interfaces provide a non-invasive and accessible alternative. Such systems eliminate the need for physical instrumentation, offering a cost-effective and hands-free setup. However, such interfaces also present implementation challenges. Monocular RGB cameras lack inherent depth information, complicating 3D pose estimation. The absence of direct depth information in monocular systems results in significantly reduced accuracy compared to RGB-D approaches, with studies showing that monocular depth estimation suffers from scale ambiguity [14, 15, 16] and drift over time, particularly affecting precision in dynamic manipulation tasks [17, 18]. Moreover, many vision solutions are limited to specific gestures for discrete actions or use complex pipelines that hinder real-time responsiveness and generalization [19, 20]. RGB-D cameras, providing direct depth measurement, simplify 3D pose estimation for manipulator control and demonstrate superior accuracy compared to monocular approaches [21, 15].
@@
In order to improve the robustness and applicability of human-machine interfaces, intelligent control systems can assist by partially automating task execution, reducing the overall complexity of the task, and lowering the user’s effort. Semi-autonomous systems are presented as an alternative to complex and counterintuitive manual systems, combining decoded user intentions with autonomous control modules. This work introduces a vision-based teleoperation shared-control framework designed to overcome the limitations outlined above, providing intuitive, real-time control of a quadruped’s manipulator. Our system uses a depth camera for robust 3D perception to perform hand pose estimation. The proposed framework tracks the operator’s wrist position for continuous end-effector control and recognizes distinct hand gestures for discrete commands, enabling natural, robust, and responsive arm teleoperation. The robot’s arm follows the operator’s wrist pose while continuously checking for collisions in order to ensure safe operation.
@@
The interface enables easy switching between manual and autonomous modes via finger counting, providing a simple and low-effort method for sharing control. In autonomous mode, the robot initiates a grasp when a closed-hand gesture is detected over a detected object. The system identifies graspable items from a predefined list, selects the one with the highest confidence and closest proximity, and provides its location to help the robot locate and pick it up. This combination of depth vision, simplified hand tracking, object recognition, and intuitive mode switching aims to create a holistic, practical, and shared-control system. Importantly, it requires neither wearable devices nor traditional interfaces, addressing cost and setup barriers and offering a low-cost, accessible alternative for complex telemanipulation, especially where other systems are impractical. Experiments demonstrated a favorable balance of intuitiveness, responsiveness, and practicality for a series of pick-and-place tasks, as shown in Fig. 1.
@@
The rest of the paper is organized as follows: Section II explains how the framework was implemented and deployed. Section III presents the apparatus used and the experiments performed to evaluate the framework’s efficiency, along with the experimental results of this study. Section IV concludes the paper.
@@ II-A Camera-Based Motion Capture System
Figure 2: Overview of the proposed control pipeline, illustrating the flow from motion capture to robot actuation. A manual teleoperation module allows the user to control the robot in an intuitive manner using an external camera. A semi-autonomous module, capable of autonomously grasping an object, can be triggered by the user by closing their hands. This framework was employed to perform a pick-and-place task using different objects.
@@ II-B Control Framework
The Spot robotic arm is controlled using the Boston Dynamics Software Development Kit (SDK) [26], ROS-based interfaces, and the MoveIt motion planning framework [27], enabling precise end-effector pose control. Motion commands are generated based on the estimated 3D wrist pose of the user, $p_{\text{marker}}^{\text{wrist}}$, acquired through the camera-based motion capture system detailed in Section II-A. These commands, representing the desired end-effector pose for the robot, are transmitted to the Spot SDK via ROS, enabling real-time control of the arm. Gripper actions, such as opening and closing for object manipulation, are similarly triggered by discrete gesture-based inputs recognized by an independent gesture recognition module, which is based on hand landmarks from MediaPipe. This shared-control architecture enables intuitive teleoperation while ensuring robust and safe object manipulation. The overall control pipeline is illustrated in Fig. 2. This hybrid architecture leverages the advantages of both the proprietary Boston Dynamics SDK, which provides robust, low-level access to Spot’s hardware, and the flexibility and extensibility of the open-source ROS ecosystem. ROS, together with MoveIt, manages bidirectional command and state translation, acting as the integration point for perception, planning, and actuation.
@@ II-B4 Implementation Details
The control loop, encompassing motion capture data processing, command generation, and communication with the robot, is developed for real-time performance. Perception tasks (object detection, if active) and motion planning are executed in parallel where feasible. The code for integrating user input, MoveIt functionalities, and dispatching commands to Spot, alongside Docker files, are publicly available in our repository (see footnote in Section II-A). Debugging overlays are provided for visualizing all perception and control stages, aiding in parameter tuning and troubleshooting.
@@ III-B Simulation Experiments
Figure 5: Decoded position of the user’s wrist using the pose estimation of the wrist and using the Optitrack system, and the translated robot’s end-effector position captured using the Optitrack system. These measurements were used to calculate the positional error, using the Optitrack’s values as ground-truth.
@@ III-D Precision analysis
In order to perform a more comprehensive analysis of the proposed framework, we attached reflexive markers to the user’s wrist and robot’s end-effector to capture its position using an Optitrack system comprised of six Primex, 41 cameras and used this as ground truth to perform an analysis of the pose estimation’s positional error, as the Optitrack system can achieve 3D accuracy of up to +/- 0.10 mm. The results are shown in Fig. 5. The positional error between the user’s wrist from the pose estimation and the actual robot’s end effector is shown in Fig. 6, alongside the user’s wrist velocity. A mean positional error of 0.07 m was achieved, and it can be observed that the error increases as the wrist velocity increases. This is expected, as the robot’s end effector takes some time to reach the goal position, leading to this increased error.
@@ IV Conclusion
This paper presents an intuitive and assistive vision-based telemanipulation framework for controlling a robotic arm system of a four-legged robot. An external camera is used to track the operator’s arm movement, offering intuitive control of the robotic system. The framework enables the user to control the end-effector position and orientation, where the position is obtained by tracking the user’s wrist position, and orientation tracking is achieved by estimating the pose of the operator’s hand. The proposed system also enables collision detection, ensuring safe operation. The framework was experimentally validated in a telemanipulation task, where two users successfully performed pick-and-place tasks in real-time. Future work will focus on implementing a path planner within the framework to achieve obstacle avoidance. Moreover, we also plan to implement a potential fields scheme [29] in order to improve the robustness of our proposed method through a shared-control approach, incorporating more participants, object diversity, and longer trials into the experiments. The proposed framework will be further validated by performing user and comparative studies with traditional interfaces.